given the points below, find xy. round to the nearest hundredth. x(-1,2) and y(4,0)


1
Expert's answer
2021-09-21T16:12:38-0400

"xy=\sqrt{(4-(-1))^2+(0-2)^2}\\xy=\sqrt{(4+1)^2+(-2)^2}\\xy=\sqrt{(5)^2+(-2)^2}\\xy=\sqrt{25+4}"

"xy= \sqrt{29}=5.38516481"

hence

"xy=5.38"